# Vertex-minimal hyperbolic origami 2-torus

### v1

@@ Abstract
In this paper, we explore geometrizations of triangulated hyperbolic 2-tori $\Sigma_{2}$ in the hyperbolic 3-space $\mathbb{H}^{3}$. We show that for all $n\geq 10$, there exists a triangulation of $\Sigma_{2}$ with $n$ vertices and an embedding $S:\Sigma_{2}\hookrightarrow\mathbb{H}^{3}$ that restricts to isometries on the triangles. We call these embeddings hyperbolic origami 2-tori. Since the lower bound $n\geq 10$ is combinatorially sharp, this paper settles the question of minimum number of vertices required to obtain a hyperbolic embedding.
@@ 1.1 Context and Problem Statement
A hyperbolic 2-torus $\Sigma_{2}$ is a quotient of $\mathbb{H}^{2}/\Gamma$ that is topologically an orientable closed surface of genus 2. Here, $\Gamma$ is a discrete subgroup of $\operatorname{Isom}(\mathbb{H}^{2})$ isomorphic to the group with presentation $\langle a,b,c,d\,|\,[a,b][c,d]\rangle$, where $a,b]=aba^{-1}b^{-1}$ denotes the commutator of the two elements, and the action of $\Gamma$ on $\mathbb{H}^{2}$ is free and properly discontinuous. Every hyperbolic 2-torus $\Sigma_{2}$ inherits the structure of a Riemannian manifold of constant sectional curvature -1 from its universal cover. A hyperbolic triangulation $\mathcal{T}$ of $\Sigma_{2}$ is a triangulation of $\Sigma_{2}$ whose edges are all geodesics. A hyperbolic origami 2-torus is a map $S:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$ that restricts to isometries on the triangles. Concretely, we think of a hyperbolic origami 2-torus $S$ as gluing hyperbolic triangles together in $\mathbb{H}^{3}$, where the cone angles at each vertex are all $2\pi$.
The study of graph embeddings and triangulations of closed surfaces can be dated back to the late 19th century [Hea90]. The minimum number of vertices $n_{0}$ to triangulate a closed surface of Euler characteristic $\chi$ is given by
$n_{0}\geq\bigg\lfloor\frac{7+\sqrt{49-24\chi}}{2}\bigg\rfloor$ (1)
Equation (1) is proved by Ringel [Rin55] for the non-orientable cases and later by Jungerman and Ringel [JR80] for the orientable cases. However, the bounds in Equation (1) is not tight for the 2-torus. In [Hun78], Huneke showed that the minimum number of vertices required to triangulate a 2-torus is 10, in contrast with a looser lower bound of 9 from Equation (1). In 2008, Lutz [Lut08] enumerated all 865 isomorphism classes of 10-vertex triangulations of the 2-torus.
Geometers were interested in whether these triangulated surfaces admit piecewise-affine embeddings into $\mathbb{R}^{3}$ that are affine on the triangle pieces [BB87, Bre87, BB89, BZ95, HLZ06]. These embeddings are called polyhedral embeddings. A polyhedral embedding of a 10-vertex triangulation of the 2-torus was first discovered by Brehm [Bre81]. Later, Hougardy, Lutz, and Zelke [HLZ07] developed a computer algorithm to enumerate polyhedral embeddings of all 10-vertex triangulations of the 2-torus. However, a foundamental problem remains unsolved: if we take into account the hyperbolic structures on a 2-torus, these polyhedral embeddings fail to preserve them. First, it’s impossible to map hyperbolic triangles isometrically to geodesic triangles in $\mathbb{R}^{3}$ due to Gauss-Bonnet theorem. Thus, it is more natural to consider embeddings of $(\Sigma_{2},\mathcal{T})$ into $\mathbb{H}^{3}$, where the images of triangles in $\mathcal{T}$ are isometric geodesic triangles in the hyperbolic 3-space.
@@ 1.2 The Candidate Model 2-Torus
For all $n\geq 10$, there exists a triangulation $\mathcal{T}$ of $\Sigma_{2}$ with $n$ vertices and a hyperbolic origami 2-torus $S:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$.
@@
The information of a hyperbolic origami 2-torus is captured entirely by the triangulation $\mathcal{T}$ and the image of the vertices of $\mathcal{T}$ under the map $S:(\Sigma_{2},\mathcal{T})\rightarrow\mathbb{H}^{3}$, as long as the cone angles of the vertices of $\mathcal{T}$ are all $2\pi$. We will show the existence of a 10-vertex triangulation $\mathcal{T}$ and a hyperbolic origami 2-torus $S$. To get a hyperbolic origami 2-torus with arbitrary number of vertices $n\geq 10$, we can add vertices to the interior of the triangles in $\mathcal{T}$ to obtain triangulations $\mathcal{T}^{\prime}$ with more vertices. The map $S:\Sigma_{2}\rightarrow\mathbb{H}^{3}$ remains a hyperbolic origami 2-torus with respect to the subdivided triangulations $\mathcal{T}^{\prime}$.
@@
The proof of Theorem 1.1 consists of two steps. First, we find an embedded hyperbolic origami 2-torus $\hat{S}$ that is “robustly embedded” and “almost flat.” Next, we show that we can get a flat one by purturbing the coordinates of $\hat{S}$. Figure 1 shows the hyperbolic polyhedral 2-torus $\hat{S}$ embedded in the Beltrami–Klein model of the hyperbolic 3-space, where the geodesics are straight lines and geodesic triangles are triangles in the Euclidean sense. In Figure 2 we attempt to draw the induced tiling of the universal cover $\mathbb{H}^{2}$ by the hyperbolic triangulation on $\mathcal{T}$ on $\Sigma_{2}$ by isometrically “lifting” the triangles in $\hat{S}$ to $\mathbb{H}^{2}$ and gluing them together. Since $\hat{S}$ is not an origami 2-torus, we can only draw the tiling in the universal cover up to some error (about $10^{-28}$), but the error is too small to be discernable by human eyes so that we can still make sense of the approximated tiling. See Figure 3, 4, 5 for the intersection of $\hat{S}$ with the $xy$-plane, xzxz-plane, and $yz$-plane.
@@
Figure 4: The intersection of $\hat{S}$ with the plane $y=0$, which is the boundary of two simply-connected components colored in yellow. The boxes contain the manified view of their respective regions in the original slice.
We explain how we chose the triangulation $\mathcal{T}$ and obtained these coordinates. We start with all triangulations $\mathcal{T}_{n}=(V_{n},E_{n},F_{n})$ and polyhedral embeddings $S_{n}:(\Sigma_{2},\mathcal{T}_{n})\hookrightarrow\mathbb{R}^{3}$ from [HLZ07] for all 865 triangulations $n=1,\ldots,865$, where $S_{n}$ maps vertices of $\mathcal{T}_{n}$ to the integer lattice $\mathbb{Z}^{3}\cap[0,4]^{3}$. First, we apply a translation and scale the vectors so that the center-of-mass of the vertices is at the origin, and that $\max_{X\in V_{n}}\|X\|=\frac{1}{2}$. This ensures that $S_{n}(\Sigma_{2})\subset B_{1}(0)$, so we can treat $S_{n}$ as an embedding into the Beltrami–Klein model of $\mathbb{H}^{3}$. Then, for each $n$, we compute the hyperbolic angles of all triangles in $S_{n}$ and all the cone angles. We found that $S_{265}$ has the minimal cone angle difference to $2\pi$ among all other embeddings. After that, we apply a hill-climbing algorithm by randomly purturbing the coordinates of $S_{265}$ to minimize the objective function
@@
Figure 5: The intersection of $\hat{S}$ with the plane $x=0$, which is the boundary of a region that is not simply-connected. Again, the boxes contain the manified view of their respective regions in the original slice.
@@
This paper is organized as follows: In $\S$2, we discuss the properties of the Beltrami–Klein model and explain how we incorporate floating-point arithmetic into the proofs. In $\S$3, we show that the cone angles of $\hat{S}$ differ from $2\pi$ by at most $10^{-28}$, and that $\hat{S}$ stays embedded after purturbing its $z$-coordinates by at most $10^{-7}$. In $\S$4, we analyze the Jacobian matrix of $\Theta(z)$ to show that there exists a flat embedded hyperbolic origami 2-torus $S$ near $\hat{S}$.
@@ 1.3 Acknowledgements
This project is supported by an N.S.F. Research Grant DMS-2505281. I would like to thank my advisor, Richard Schwartz, for introducing this problem to me and kindly explaining to me his proof of the minimal-vertex paper torus problem. Much of the methodology used in this paper is inspired by [Sch25]. I would like to thank Alba Málaga, Fabian Lander, Stepan Paul, and Steve Trettel for helpful discussions on this topic. I would like to thank Frank Lutz for his manifolds webpage and Stefan Hougardy, Frank H. Lutz, and Mariano Zelke for keeping a database of piecewise-affine embedded 2-tori with vertex-minimal triangulations, which helped me find prototypes for the cadidate 2-torus in this paper.
@@ 2.1 The Beltrami–Klein Model
where $\langle\cdot,\cdot\rangle$ denotes the Euclidean dot product. We will write $\langle V,W\rangle_{X}=g|_{X}(V,W)$ and $\|V\|_{X}=\sqrt{g|_{X}(V,V)}$ the norm of $V\in T_{X}\mathbb{H}^{n}$.
Since geodesics in $\mathbb{H}^{3}$ are straight lines, a totally geodesic triangle in $\mathbb{H}^{3}$ with vertices $(X,Y,Z)$ is the same as the Euclidean triangle in the unit ball. We may compute the hyperbolic angle $\theta$ at $X$ as follows: Let V=YXV=Y-X and W=YXW=Y-X. We may identify $V$ and $W$ with velocity vectors in $T_{X}\mathbb{H}^{3}$ of the geodesic from $X$ to $Y$ and $X$ to $Z$ respectively. Consider the following function:
@@ 3 The Candidate 2-Torus
In this section, we show that the cone angles of our candidate 2-torus $\hat{S}$ are very close to $2\pi$. We then show that if we apply purturbation to the $z$-coordinates of $\hat{S}$, the resulting 2-torus stays embedded as long as the purturbation is small enough.

### v2

@@ Abstract
We show that there exists a geodesic triangulation $\mathcal{T}$ of a hyperbolic genus 2 surface $\Sigma_{2}$ with 10 vertices and an isometric polyhedral embedding $S:\Sigma_{2}\hookrightarrow\mathbb{H}^{3}$ that sends the triangles in $\mathcal{T}$ to geodesic triangles in $\mathbb{H}^{3}$. We call this type of embedding a hyperbolic origami 2-torus. Since 10 is the combinatorially minimum number of vertices required to triangulate a genus 2 surface, this paper settles the question of minimum number of vertices required to obtain a hyperbolic origami 2-torus.
@@ 1.1 Context and Problem Statement
A hyperbolic 2-torus $\Sigma_{2}$ is a quotient of $\mathbb{H}^{2}$ by the action of a surface group $\Gamma<\operatorname{Isom}^{+}(\mathbb{H}^{2})$ such that $\mathbb{H}^{2}/\Gamma$ is a genus 2 surface. Here, $\Gamma$ is isomorphic to the group with presentation $\langle a,b,c,d\,|\,[a,b][c,d]\rangle$, where $a,b]=aba^{-1}b^{-1}$ denotes the commutator of the two elements, and the action of $\Gamma$ on $\mathbb{H}^{2}$ is free and properly discontinuous. Every hyperbolic 2-torus $\Sigma_{2}$ inherits the structure of a Riemannian manifold of constant sectional curvature -1 from $\mathbb{H}^{2}$. A hyperbolic origami 2-torus is a map $S:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$, where $\mathcal{T}$ is a geodesic triangulation of $\Sigma_{2}$, and $S$ restricts to isometries on the triangles of $\mathcal{T}$. Concretely, we think of a hyperbolic origami 2-torus $S$ as gluing hyperbolic triangles together in $\mathbb{H}^{3}$, where the cone angles at each vertex are $2\pi$.
The study of graph embeddings and triangulations of closed surfaces can be dated back to the late 19th century [Hea90]. The minimum number of vertices $n_{0}$ to triangulate a closed surface of Euler characteristic $e$ is given by
$n_{0}\geq\bigg\lfloor\frac{7+\sqrt{49-24e}}{2}\bigg\rfloor$ (1)
Equation (1) is proved by Ringel [Rin55] for the non-orientable cases and later by Jungerman and Ringel [JR80] for the orientable cases. However, the bound in Equation (1) is not tight for the 2-torus. In [Hun78], Huneke showed that the minimum number of vertices required to triangulate a 2-torus is 10, in contrast with a looser lower bound of 8 from Equation (1). In 2008, Lutz [Lut08] enumerated all 865 isomorphism classes of 10-vertex triangulations of the 2-torus.
Geometers were interested in whether these triangulated surfaces admit piecewise-affine embeddings into $\mathbb{R}^{3}$ that are affine on the triangle pieces [BB87, Bre87, BB89, BZ95, HLZ06]. These embeddings are called polyhedral embeddings. A polyhedral embedding of a 10-vertex triangulation of the 2-torus was first discovered by Brehm [Bre81]. Later, Hougardy, Lutz, and Zelke [HLZ07] developed a computer algorithm to enumerate polyhedral embeddings of all 10-vertex triangulations of the 2-torus. However, a fundamental problem remains unsolved: if we take into account the hyperbolic structures on a 2-torus, these polyhedral embeddings fail to preserve them. First, it’s impossible to map hyperbolic triangles isometrically to geodesic triangles in $\mathbb{R}^{3}$ due to the Gauss–Bonnet theorem. Thus, it is more natural to consider embeddings of $(\Sigma_{2},\mathcal{T})$ into $\mathbb{H}^{3}$, where the images of triangles in $\mathcal{T}$ are isometric geodesic triangles in the hyperbolic 3-space.
@@ 1.2 The Candidate Model 2-Torus
There exists a hyperbolic 2-torus $\Sigma_{2}$ along with a 10-vertex geodesic triangulation $\mathcal{T}$ and a hyperbolic origami 2-torus $S:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$.
@@
The information of a hyperbolic origami 2-torus is captured entirely by the triangulation $\mathcal{T}$ and the image of the vertices of $\mathcal{T}$ under the map $S:(\Sigma_{2},\mathcal{T})\rightarrow\mathbb{H}^{3}$, as long as the cone angles of the vertices of $\mathcal{T}$ are all $2\pi$. We remark that Theorem 1.1 implies the existence of hyperbolic origami 2-tori with $n$ vertices for all $n\geq 10$; one simply uses the 10-vertex embedding and adds vertices to the interior of the triangles in $\mathcal{T}$ to obtain triangulations with more vertices.
@@
The proof of Theorem 1.1 consists of two steps. First, we find a polyhedral 2-torus $\hat{S}$ that is “robustly embedded” and “almost isometric.” Next, we argue that an isometric one exists by perturbing the coordinates of $\hat{S}$. Figure 1 shows the polyhedral 2-torus $\hat{S}$ embedded in the Beltrami–Klein model of the hyperbolic 3-space (recall that in the Klein model, geodesics are straight lines and geodesic triangles are Euclidean triangles). In Figure 2, we draw the induced tiling of the universal cover $\mathbb{H}^{2}$ by the “almost geodesic” triangulation on $\Sigma_{2}$ by “almost isometrically” lifting the triangles in $\hat{S}$ to $\mathbb{H}^{2}$ and gluing them together. Since $\hat{S}$ is not a hyperbolic origami 2-torus, the picture of the tiling in the universal cover is only meant to be an approximation up to some error (about $10^{-28}$), but the error is too small to be discernible by human eyes so that visually the picture looks reasonable. See Figure 3, 4, 5 for the intersection of $\hat{S}$ with the $xy$-plane, xzxz-plane, and $yz$-plane.
@@
Figure 4: The intersection of $\hat{S}$ with the plane $y=0$, which is the boundary of two simply-connected components colored in yellow. The boxes contain the magnified view of their respective regions in the original slice.
We explain how we chose the triangulation $\mathcal{T}$ and obtained these coordinates. We start with all triangulations $\mathcal{T}_{n}=(V_{n},E_{n},F_{n})$ and polyhedral embeddings $S_{n}:(\Sigma_{2},\mathcal{T}_{n})\hookrightarrow\mathbb{R}^{3}$ from [HLZ07] for all 865 triangulations $n=1,\ldots,865$, where $S_{n}$ maps vertices of $\mathcal{T}_{n}$ to the integer lattice $\mathbb{Z}^{3}\cap[0,4]^{3}$. First, we apply a translation and scale the vectors so that the center-of-mass of the vertices is at the origin, and that $\max_{X\in V_{n}}\|X\|=\frac{1}{2}$. This ensures that $S_{n}(\Sigma_{2})\subset B_{1}(0)$, so we can treat $S_{n}$ as an embedding into the Klein model of $\mathbb{H}^{3}$. Then, for each $n$, we compute the hyperbolic angles of all triangles in $S_{n}$ and all the cone angles. We found that $S_{265}$ has the minimal cone angle difference to $2\pi$ among all other embeddings. After that, we apply a hill-climbing algorithm by randomly perturbing the coordinates of $S_{265}$ to minimize the objective function
@@
Figure 5: The intersection of $\hat{S}$ with the plane $x=0$, which is the boundary of a region that is not simply-connected. Again, the boxes contain the magnified view of their respective regions in the original slice.
1.3 A 12-Vertex Model with 2-Fold Symmetry
The candidate model $\hat{S}$ is rather unsatisfactory in the sense that it doesn’t admit any symmetry. On the other hand, the 8-vertex paper torus model from [Sch25] has a 2-fold symmetry that corresponds to a hyperelliptic involution, where the four fixed points are midpoints of 4 edges. Since every flat torus admits a hyperelliptic involution, Schwartz’s construction effectively reduced the dimension of the search space to find a family of paper torus in the hope of realizing every flat structure. Very recently, Doyle and Schwartz [DS25] found a family of 8-vertex flat polyhedral tori that realize any flat torus without reflection symmetry, which constitutes a dense subset in the moduli space of flat tori.
Figure 6: The embedding $\hat{S}_{12}$ with 2-fold rotational symmetry about the $z$-axis. Left: front view from the positive $x$-direction. Middle: bird’s-eye view from the positive $z$-direction. Right: view from the negative $z$-direction.
Motivated by this finding, I searched for a hyperbolic origami 2-torus with 2-fold symmetry that corresponds to a hyperelliptic involution. One first observes that if the order-2 symmetry acts freely on the set of vertices and fixes 6 points, then it must fix midpoints of 6 edges, which requires at least 12 vertices. I started by finding piecewise-linear embeddings of a 12-vertex triangulation of the 2-torus along with a candidate model with 2-fold symmetry into $\mathbb{R}^{3}$. A natural starting point is to consider the 7-regular triangulations, which has the largest symmetry group. I first checked whether the 7-regular triangulations of the 2-torus from [DU06] admit piecewise-linear embeddings into $\mathbb{R}^{3}$ with 2-fold symmetry. Then, I applied the same procedure as before to numerically approximate a hyperbolic origami 2-torus with 12 vertices and 2-fold symmetry. This led me to the following 12-vertex triangulation $\mathcal{T}$ with vertices indexed by $\{0,\ldots,11\}$. The indices of the 28 oriented triangles in $\mathcal{T}$ are listed below:
$\begin{matrix}(0,1,5)&(0,4,8)&(0,5,11)&(0,6,1)&(0,7,6)&(0,8,7)&(0,11,4)\\ (1,2,7)&(1,6,2)&(1,7,8)&(1,8,9)&(1,9,5)&(2,3,7)&(2,6,10)\\ (2,8,3)&(2,9,8)&(2,10,9)&(3,4,9)&(3,8,4)&(3,9,10)&(3,10,11)\\ (3,11,7)&(4,5,9)&(4,10,5)&(4,11,10)&(5,6,11)&(5,10,6)&(6,7,11)\end{matrix}$ (5)
Notice that $\mathcal{T}$ admits a $\mathbb{Z}/12\mathbb{Z}$ symmetry generated by the map $\sigma:i\mapsto(i+1)\mod 12$. The hyperelliptic involution is given by $\sigma^{6}$. Using this triangulation, I found a polyhedral embedding $\hat{S}_{12}:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$ that satisfies $R\circ\hat{S}_{12}=\hat{S}_{12}$, where $R\in\operatorname{Isom}^{+}(\mathbb{H}^{3})$ is a rotation by $\pi$ about the $z$-axis. I checked that the cone angles of $\hat{S}_{12}$ are bounded away from $2\pi$ by $10^{-10}$. I provide the coordinates of $\hat{S}_{12}$ in $\S$5.3.
Figure 6 contains the visualization of the embedding $\hat{S}_{12}$ in the Klein model. It is very curious that when we use the Klein model to visualize $\hat{S}_{12}$, its appearance resembles a “double-deck pup tent” (the pup tent refers to the paper tori construction by Schwartz in [Sch25]). This hints at the possibility to generalize this construction.
Figure 7: The intersection of the 12-vertex model with the plane $x=0$ (left), $y=0$ (middle), and z=0.05z=-0.05 (right). One can clearly see the six midpoints of edges $(i,i+6)$ that are fixed by the rotational symmetry from the $x=0$ and $y=0$ slices.
The midpoints of the six edges $(i,i+6)$ all lie on the $z$-axis, which can be seen by slicing the embedding with planes that contains the $z$-axis. In Figure 7, we show the intersection of $\hat{S}_{12}$ with the planes $x=0$, $y=0$, and z=0.05z=-0.05. One can clearly see the six fixed points of the hyperelliptic involution from the $x=0$ and $y=0$ slices. Finally, in Figure 8, we show the induced triangle tiling of the universal cover $\mathbb{H}^{2}$ from $\hat{S}_{12}$ in the Poincaré disk model. In this visualization, the pairs of triangles $(i,j,k)$ and $(i+6,j+6,k+6)$ are colored the same way to reflect the 2-fold symmetry.
Figure 8: The isometrically induced 7-regular triangle tiling in the universal cover $\mathbb{H}^{2}$, visualized in the Poincaré disk model. The vertices $z_{0}$ and $z_{6}$ are fixed on the real axis, with $\Re(z_{0})<0$ and $z_{6}=-z_{0}$.
1.4 Paper Organization
@@
This paper is organized as follows: In $\S$2, we remind the readers of the properties of the Klein model and explain how we incorporate floating-point arithmetic into the proofs. In $\S$3, we show that the cone angles of $\hat{S}$ differ from $2\pi$ by at most $10^{-28}$, and that $\hat{S}$ stays embedded after perturbing its $z$-coordinates by at most $10^{-7}$. In $\S$4, we analyze the Jacobian matrix of $\Theta(z)$ to show that there exists a hyperbolic origami 2-torus $S$ near $\hat{S}$.
An ongoing project is to generalize the 12-vertex construction to a family of hyperbolic origami 2-tori and check which hyperbolic structures can be realized.
1.5 Accompanying Programs
I have written programs in both Python and Mathematica to visualize the candidate models and slicing them with planes. Readers can access them via the following link:
@@ 1.6 Acknowledgements
This project is supported by an N.S.F. Research Grant DMS-2505281. I would like to thank my advisor, Richard Schwartz, for inspiring me to work on this problem. Much of the methodology used in this paper is inspired by [Sch25]. I would like to thank Peter Doyle, Fabian Lander, Alba Málaga, Stepan Paul, Steve Trettel, and Bena Tshishiku for helpful discussions on this topic. I would like to thank Frank Lutz for his manifolds webpage [Lut17] and Stefan Hougardy, Frank H. Lutz, Mariano Zelke for keeping a database of piecewise-affine embedded 2-tori with vertex-minimal triangulations, which helped me find prototypes for the candidate 2-torus in this paper.
@@ 2.1 The Beltrami–Klein Model
where $\langle\cdot,\cdot\rangle$ denotes the Euclidean dot product. We will write $\langle V,W\rangle_{X}=g|_{X}(V,W)$ and $\|V\|_{X}=\sqrt{g|_{X}(V,V)}$ for the norm of $V\in T_{X}\mathbb{H}^{n}$.
Since geodesics in $\mathbb{H}^{3}$ are straight lines, a totally geodesic triangle in $\mathbb{H}^{3}$ with vertices $(X,Y,Z)$ is the same as the Euclidean triangle in the unit ball. We may compute the hyperbolic angle $\theta$ at $X$ as follows: Let V=YXV=Y-X and W=ZXW=Z-X. We may identify $V$ and $W$ with velocity vectors in $T_{X}\mathbb{H}^{3}$ of the geodesic from $X$ to $Y$ and $X$ to $Z$ respectively. Consider the following function:
@@ 3 The Candidate 2-Torus
In this section, we show that the cone angles of our candidate 2-torus $\hat{S}$ are very close to $2\pi$. We then show that if we apply perturbation to the $z$-coordinates of $\hat{S}$, the resulting 2-torus stays embedded as long as the perturbation is small enough.